OpenStudy (anonymous):

Find the absolute, relative and percentage errors when true value is 0.50 and its calculated value was 0.49.

OpenStudy (ybarrap):

Absolute: $$ \left | .49-.50\right |$$ Relative: $$.49-.50$$ Percentage $$ \cfrac{.49-.50}{.50}\times 100 $$ Make sense?

OpenStudy (anonymous):

So the absolute error is 0.01, the relative error is -0.01, and the percentage error is -1%?

OpenStudy (ybarrap):

check your calculation for the last one

OpenStudy (anonymous):

Is it really -2%?

OpenStudy (ybarrap):

yes

OpenStudy (anonymous):

But for other problems, if I got 0.0005 as the relative error, then people say that the percentage error should be 0.05%.

OpenStudy (ybarrap):

So using this as a definition of relative and percentage error: http://en.wikipedia.org/wiki/Relative_error, you would be correct. Relative Error: $$ \left |\cfrac{.49-.50}{.50}\right | $$ Percentage Error=Relative Error \(\times 100\)